OpenStudy (anonymous):

solve: log[12](x+1)-log[12](4)=log[12](16). x=3 x=12 x=63 x=64

OpenStudy (mertsj):

\[\log_{12}\frac{x+1}{4}=\log_{12}16 \] (x+1)/4=16

OpenStudy (anonymous):

thanks